Please help me solve this equation
6×+5=3×-10

Found 2 solutions by rfer, richwmiller:
Answer by rfer(16322)

6x-3x=-10-5
3x=-15
x=-15/3
x=-5

Answer by richwmiller(17219)

What is the problem?